# Pathways on the Image Manifold: Image Editing via Video Generation

### v1

@@ 6 Limitations
While our approach addresses key shortcomings in current image editing methods, it also introduces unique challenges. For example, natural camera motion sometimes appears in video sequences, which, when replicated in generated content, can lead to unintended perspective shifts. As with all generative models, video models are trained on specific data domains, making it challenging to produce results that deviate significantly from the model’s training data, which predominantly includes real-world transformations. Despite these challenges, F2F demonstrates success in several cases. For example, in the middle row of Figure 1, the vase is ”magically” filled with green water, showcasing the model’s ability to perform imaginative edits. Additionally, our method is computationally intensive, as transforming an image into a video sequence is resource-heavy and often slower than other image editing methods. However, video generation efficiency is advancing rapidly, as demonstrated by models like Runway Turbo [^2] Runway Turbo API and LTX-Video [^3] LTX-Video, which can execute the process in seconds, potentially making this approach significantly less resource-intensive in the future.
@@ 7 Conclusions
Our work opens several promising directions. First, fine-tuning video generators specifically for image editing presents an exciting opportunity, including straightforward solutions like enforcing static camera scenarios or using datasets curated for editing tasks, alongside more complex approaches yet to emerge. Secondly, a key direction could involve reducing the overhead of full video generation while preserving the benefits of gradual, temporally coherent transformations, potentially enhancing both the efficiency and speed of editing.
@@ Appendix A Additional Editing Examples
We provide additional qualitative examples of the editing results on the TEdBench benchmark, generated based on the experimental setup detailed in Section 5.2. These examples further supplement those presented in Figure 1 and Figure 4. Each example is accompanied by the temporal editing caption used to perform the edit, which was generated using the method described in Section 3.1.
@@ B.1 VLM Instruction
’Write a one-sentence description of a short video that begins with the provided image and smoothly transitions into a scene of a ”CAPTION”, highlighting how elements in the image undergo changes or movement over time. Keep the description simple, concise and short, focusing only on essential changes and actions without altering unnecessary details. Avoid mentioning elements that do not contribute to the main change needed, and focus the description on the main transitions. Do not add objects that are not in the original image or described in the final scene. The camera should remain static unless movement is absolutely necessary. Ensure all transitions happen within a few second duration without mentioning the length or using the word ”video”.’
Here, ”CAPTION” is replaced with the target caption specific to the image. Additionally, as explained, in-context learning is employed to provide the VLM with examples alongside the instruction. Before processing the desired source image and edit prompt, the instruction is presented to the VLM nine times, each paired with a distinct example consisting of a source image, target caption, and corresponding temporal editing caption. Examples of these are illustrated in Figure S9.
@@ C.1 VLM Instruction
‘The image displays the source photo at the top, with a collage of 12 edited versions beneath it. The target edit image caption was: “CAPTION”. Your task is to choose the image from 1 to 12 that best follows this edit fully and naturally. If none of the images follows the edit, select image 0. If multiple images follow the edit equally, prioritize the one with the lowest number possible. Avoid selecting images that appear to follow the edit but are not edits of the original image. Additionally, avoid images where camera motion, zoom, or image quality differs significantly, or where the content does not appear stable relative to the original source. Respond with: “The selected edit is:x” where x is the number of your chosen edit.’
Here, ”CAPTION” refers to the target editing caption. Examples of the collages can be found in Figure S10.
Figure S10: Frame Selection Collage. The target editing caption for this example is: “A photo of a cat yawning.”.
@@ Appendix D Editing Manifold Pathway
Appendix E PosEdit
In Section 5.3, we outline the construction of a human pose editing dataset. This dataset encompasses 58 editing tasks, distributed across 8 distinct action categories, featuring 8 different subjects. The source images consistently depict a neutral standing pose with arms relaxed at the sides, while the target poses vary according to the edit category. Each editing category is paired with a target caption and a temporal caption. Figure S12 and Figure S13 illustrates examples for each action category.
@@ Appendix E PosEdit
Appendix F Human Survey
As detailed in Section 5.4, we conducted a human evaluation survey to assess our method’s performance based on real user preferences. Following the framework in [47], the survey questions evaluated (1) the accuracy of the edit relative to the prompt and (2) the quality of the edit, defined as the preservation of visual fidelity to the source image. Each participant reviewed 20 edits, comparing our method with LEDITS++. Examples of the pages shown to the evaluators are provided in Figures S14 and S15.
Figure S14: Survey Example
Figure S15: Survey Example
Appendix G Additional Vision Tasks Captions

### v4

@@ 6 Limitations
While our approach addresses key shortcomings in current image editing methods, it also introduces unique challenges. For example, natural camera motion sometimes appears in video sequences, which, when replicated in generated content, can lead to unintended perspective shifts. As with all generative models, video models are trained on specific data domains, making it challenging to produce results that deviate significantly from the model’s training data, which predominantly includes real-world transformations. Despite these challenges, F2F demonstrates success in several cases. For example, in the middle row of Figure 1, the vase is "magically" filled with green water, showcasing the model’s ability to perform imaginative edits, with additional examples discussed in Section H of the supplementary. Additionally, our method is computationally intensive, as transforming an image into a video sequence is resource-heavy and often slower than other image editing methods. However, video generation efficiency is advancing rapidly, as demonstrated by models like Runway Turbo [^2] Runway Turbo API and LTX-Video [17], which can execute the process in seconds, making this approach increasingly less resource-intensive. Additionally, optimizing the number of frames per edit—currently fixed at 49 for CogVideoX—could enable faster editing possibilities.
@@ 7 Conclusions
Our work opens several promising directions. First, fine-tuning video generators specifically for image editing presents an exciting opportunity, including straightforward solutions like enforcing static camera scenarios or using datasets curated for editing tasks, alongside more complex approaches yet to emerge. Secondly, a key direction could involve reducing the overhead of full video generation while preserving the benefits of gradual, temporally coherent transformations, potentially enhancing both the efficiency and speed of editing. The speed can be affected both in terms of per-frame average generation speed or the number of frames generated per edit, which is currently fixed at 49 for CogVideoX.
@@ Appendix A Additional Editing Examples
We provide additional qualitative examples of the editing results on the TEdBench benchmark, generated based on the experimental setup detailed in Section 5.2. These examples further supplement those presented in Figures 1 and 4. Each example is accompanied by the temporal editing caption used to perform the edit, which was generated using the method described in Section 3.1. Furthermore, in Figure S9, we supplement Figure 1 with additional video examples generated by our method, illustrating the transition from the source image (left) to the target edit (right).
@@
Figure S9: Generated Video Examples. Additional video sequences illustrating the temporal evolution from the source image to the target edit.
@@ B.1 VLM Instruction
’Write a one-sentence description of a short video that begins with the provided image and smoothly transitions into a scene of a "CAPTION", highlighting how elements in the image undergo changes or movement over time. Keep the description simple, concise and short, focusing only on essential changes and actions without altering unnecessary details. Avoid mentioning elements that do not contribute to the main change needed, and focus the description on the main transitions. Do not add objects that are not in the original image or described in the final scene. The camera should remain static unless movement is absolutely necessary. Ensure all transitions happen within a few second duration without mentioning the length or using the word "video".’
Here, "CAPTION" is replaced with the target caption specific to the image. Additionally, as explained, in-context learning is employed to provide the VLM with examples alongside the instruction. Before processing the desired source image and edit prompt, the instruction is presented to the VLM nine times, each paired with a distinct example consisting of a source image, target caption, and corresponding temporal editing caption. Examples of these are illustrated in Figure S10.
@@ C.1 VLM Instruction
‘The image displays the source photo at the top, with a collage of 12 edited versions beneath it. The target edit image caption was: “CAPTION". Your task is to choose the image from 1 to 12 that best follows this edit fully and naturally. If none of the images follows the edit, select image 0. If multiple images follow the edit equally, prioritize the one with the lowest number possible. Avoid selecting images that appear to follow the edit but are not edits of the original image. Additionally, avoid images where camera motion, zoom, or image quality differs significantly, or where the content does not appear stable relative to the original source. Respond with: “The selected edit is:x" where x is the number of your chosen edit.’
Here, "CAPTION" refers to the target editing caption. Examples of the collages can be found in Figure S11.
Figure S11: Frame Selection Collage. The target editing caption for this example is: “A photo of a cat yawning.".
@@ Appendix D Editing Manifold Pathway
Appendix E Inference Hyperparameters
We use the 5B-parameter image-to-video version of CogVideoX, ensuring consistency by applying the same inference hyperparameters across all experiments. The model generates a fixed 49 frames per sample. During inference, we use the default denoising scheduler in CogVideoX, which is based on DDIM with V-prediction. We perform 50 denoising steps and set the classifier-free guidance scale for text conditioning to 6.0.
Appendix F PosEdit
In Section 5.3, we outline the construction of a human pose editing dataset. This dataset encompasses 58 editing tasks, distributed across 8 distinct action categories, featuring 8 different subjects. The source images consistently depict a neutral standing pose with arms relaxed at the sides, while the target poses vary according to the edit category. Each editing category is paired with a target caption and a temporal caption. Figure S13 and Figure S14 illustrates examples for each action category. Additionally, in Figure S15, we complement Figure 5 with quantitative results, demonstrating how the numerical evaluation aligns with its intended measurement objectives.
@@ Appendix F PosEdit
Figure S15: PosEdit Quantitative Evaluation Examples.
Appendix G Human Survey
As detailed in Section 5.4, we conducted a human evaluation survey to assess our method’s performance based on real user preferences. Following the framework in [49], the survey questions evaluated (1) the accuracy of the edit relative to the prompt and (2) the quality of the edit, defined as the preservation of visual fidelity to the source image. Each participant reviewed 20 edits, comparing our method with LEDITS++. Examples of the pages shown to the evaluators are provided in Figures S16 and S17.
Figure S16: Survey Example.
Figure S17: Survey Example.
Appendix H Further Ablations
We present additional ablation results demonstrating the ability of our method to handle two distinct editing challenges: (1) diverse backgrounds and (2) out-of-video-distribution edits.
Backgrounds. Similar to our image generation approach in Section 4, we generate eight images of two objects with varying backgrounds. Each image generation prompt follows one of the two templates below, where LOCATION is replaced with one of the following settings: colorful amusement park, sandy beach, magical fairy tale forest, futuristic cityscape, old library, snowy mountain peak, bustling train station, and quaint village square.
  1. 1.
    “A small brown teddy bear sitting in LOCATION.”
  2. 2.
    “A tennis ball lies in LOCATION.”
Out-of-video-distribution editing. As discussed in Section 6, one might assume that our method would fail to perform edits requiring temporal transformations that deviate significantly from typical real-world videos (the model’s training set). To show that this is not necessarily the case, we selected two unreal editing processes, using the following temporal captions:
  1. 1.
    “The teddy bear slowly rips, revealing stuffing coming out.”
  2. 2.
    “The tennis ball gradually transforms into a ripe red tomato.”
These edits depict unrealistic events that do not ordinarily occur in real-life footage. However, as shown in Fig. S18, our method successfully handles both the background variations and the out-of-distribution nature of these transformations. Moreover, the full temporal sequences in Fig. S9 demonstrate that these edits occur seamlessly without any external interaction (the teddy bear rips apart spontaneously, and the tennis ball morphs magically into a tomato).
Figure S18: Ablation Examples.
Appendix I Additional Vision Tasks Captions